# Sublogarithmic Distributed Vertex Coloring with Optimal Number of Colors

### 30 Mar 2026 v1

@@ Appendix B Concentration Inequalities
Let $S\subseteq V$ be a random subset formed by sampling each $v\in V$ independently with probability $p$. Let $\mathcal{A}$ be any collection of subsets of $V$. For each $A\in\mathcal{A}$, let $I_{A}:=\bm{1}(A\subseteq S)$ be an indicator variable for the event that $A$ is contained in $S$. Let $f:=\sum_{A\in\mathcal{A}}I_{A}$ and $\mu:=\operatorname{\mathbb{E}}[f]$. Define
@@
    for each $s\geqslant 0$, if $X\geqslant s$ then there exists a set of at most $rs$ choices whose outcomes certify that $X\geqslant s$.
@@ Appendix C Slack Generation
The functions $T$ and $D$ are $r$-certifiable with $r=2$ and $r=3$, respectively. See the appendix and Lemma B.4 for the definition of an $r$-certifiable function. $T$ and $D$ are both $2$-Lipschitz: whether a node is activated, and which color it picks affects the outcome by at most $c=2$. We apply Lemma B.4 with $b=\mathbb{E}[Z]/10-60c\sqrt{r\cdot\mathbb{E}[T]}$ (which is non-negative for $\overline{m}/q\geqslant c_{0}$ and $c_{0}$ a large enough constant):

### 01 Apr 2026 v2

@@ Appendix B Concentration Inequalities
Let $S\subseteq V$ be a random subset formed by sampling each $v\in V$ independently with probability pp. Let $\mathcal{A}$ be any collection of subsets of $V$. For each $A\in\mathcal{A}$, let $I_{A}:=\bm{1}(A\subseteq S)$ be an indicator variable for the event that AA is contained in $S$. Let $f:=\sum_{A\in\mathcal{A}}I_{A}$ and $\mu:=\operatorname{\mathbb{E}}[f]$. Define
@@
    for each $s\geqslant 0$, if $X\geqslant s$ then there exists a set of at most rsrs choices whose outcomes certify that $X\geqslant s$.
@@ Appendix C Slack Generation
The functions $T$ and $D$ are $r$-certifiable with $r=2$ and $r=3$, respectively. See the appendix and Lemma B.4 for the definition of an $r$-certifiable function. $T$ and $D$ are both 22-Lipschitz: whether a node is activated, and which color it picks affects the outcome by at most $c=2$. We apply Lemma B.4 with $b=\mathbb{E}[Z]/10-60c\sqrt{r\cdot\mathbb{E}[T]}$ (which is non-negative for $\overline{m}/q\geqslant c_{0}$ and $c_{0}$ a large enough constant):